what is the magnetic moment of disc?

Grade:12

1 Answers

Sourabh Singh IIT Patna
9 years ago
Hiii


I'll accept the non-conducting breadth

of charge:

dQ = σ dA

dQ = σ (2 π r dr)

Period of revolution:

T = 2 π / ω

differential current:

di = dQ / Δt = dQ/T

di = σ (2 π r dr) * (ω/ 2 π)

di = σω rdr

differential alluring moment:

dm = di A = di π r^2

m = ∫ di [π r^2]

m = σωπ ∫ r^3 dr [0,r]

magnetic moment vector:

m = σωπ r^4 / 4

Where m takes the administration

of the breadth agent or the angular

velocity vector, ω